OpenStudy (anonymous):

What is the slope and y-intercept of the graph of 3x + 2y = 12?

OpenStudy (anonymous):

Change this in the form y= mx + c m = slope c = y intercept

OpenStudy (anonymous):



OpenStudy (anonymous):

go from \(Ax+By=C\) to \(y=mx+b\) to do that all you have to do is get \(y\) alone on one side:\[3x+2y=12\]subtract the \(3x\):\[2y=-3x+12\]divide both sides by \(2\):\[y=-{3\over2}+6\]so from that we can conclude that the slope is \(m=\large-{3\over2}\) and the y intercept is \(b=6\) hope this helps! :) @Hiamsam

OpenStudy (anonymous):

Thanks guys!

OpenStudy (anonymous):

One more question.@yummydum

OpenStudy (anonymous):

Can you help me on this? http://gyazo.com/7f093791002c8c33808848ad12b0db10 (screenshot)

OpenStudy (anonymous):

k sure

OpenStudy (anonymous):

\(\huge{6x+24\over6}\) \(\implies\) \(\huge{{6x\over6}+{24\over6}}\) so now we just simplify and get \(\implies~~~~\huge{x+4}\)

OpenStudy (anonymous):

Thanks! again! XD

OpenStudy (anonymous):

no problem man!

OpenStudy (anonymous):

{(96, 80), (114, 95), (132, 80), (96, 95), (114, 120)} What is the difference between the values of the least and greatest elements in the domain of this relation?

OpenStudy (anonymous):

im not sure about this one but i guess you just subtract the lowest value from the highest value:\[~~~(114~,~120)\]\[\underline{-~(96~~,~~80)}\]\[~~~~(18~~,~~40)\] so i guess the answer would be \(\large(18,40)\)

OpenStudy (anonymous):

18 or 40? It only has one choice :x

OpenStudy (anonymous):

OH wait this is domain....so that means the lowest x value subtracted from the highest x value:\[\large132-96=36\]

OpenStudy (anonymous):

thanks =p

OpenStudy (anonymous):

whenever it says \(\text{domain}\) its talking about \(x\) and whenever it says \(\text{range}\) its talking about \(y\)

OpenStudy (anonymous):

Got it!

OpenStudy (anonymous):

cool :)

OpenStudy (anonymous):

Just to check is this right? http://gyazo.com/c424840d6aa38bb7e40cb20fe660fab3

OpenStudy (anonymous):

\[3x-4y=12\]subtract \(3x\):\[-4y=-3x+12\]divide by \(-4\):\[y={3\over4}x-3\]so to find the x-intercet we put \(0\) in for \(y\):\[3x-4(0)=12\]simplify:\[3x=12\]divide by \(3\):\[x=4\] so the intercept is \((4,0)\) and the y intercept is \((0,-3)\) Option \(\text{B}\)

OpenStudy (anonymous):

Thanks, again! Lol. :)

OpenStudy (anonymous):

yup